OpenStudy (anonymous):

Suppose we have a set {1,.....,r} where r >= 1, and we select three numbers independently and randomly from the set. Find the probability that a) the 3 nos are different (when r >= 3) b) the 3 nos are same (when r >= 1) c) exactly 2 nos are same (when r >= 2)

OpenStudy (anonymous):

I need help on this asap

OpenStudy (anonymous):

a) you pick some number, and it is unique with probability 1, then second number different from first has probability \(\frac{r-1}{r}\) and third number different from first two has probability \(\frac{r-2}{r}\) so the probability that all three are distinct is \(\frac{r-1}{r}\times \frac{r-2}{r}\)